Jan sold her house on December 31 and took a SS 000 mortgage as part of the payment. Thc 10-year mortgage has a 10% nominal interest rate but it cals for semiannual payments beginnng next June 30. Ne t year Jan must report on Schedule B of her IRS Form 1040 the amount of interest that was induded in the two payments she received during the year s. What is the dollar amount of each payment Jan receives Round your answer to the nearest cent How much interest was included in the first payment2 Round your answer to the nearest cent How much repayment of principal was included? Round your answer to the nearest cent How do these values change for the second payment? Select- L The portion of the payment that is applied to interest declines, while the portion of the payment that is applied to principal increases. IL. The portion of the payment that is applied to interest increases, while the portion of the payment that is applied to principal decreases. IIL. The portion of the payment that is applied to interest and the portion of the payment that is applied to principal remains the same throughout the life of the loan. V. The portion of the payment that is applied to interest declines, while the portion of the payment that is applied to principal also dedines V. The portion of the payment that is applied to interest increases, while the portion of the payment that is applied to principal also increeses b. How much interest must Jan report on Schedule B for the first year? Round youranswer to the nearest cent. d. If the payments are constant, why does the amount of interest income change over time? As the loan 1· . As amortized (pad off), the beginning balance, hence the interest charge, increases and the repayment of princpal increases. the loan is amortized (paid off), the beginning balance, hence the interest charge, dedines and the repayment of principel increases As the loan is armortized (ped off. the beginning balance, hence the interest charge, dedin 111 and the repayment of priopel declines. . As the loan is arnortined (paid off. the beginning balance, hence tho interest charge, increases and the repayment of principal declines. V. As the loan is amortized (paid off), the beginning balarke declines, but the interest charge and the repayment of principal remain the same.
Answer #1

As per rules I am answering the first 4 subparts of the question

A:

We first need to find the instalment amounts

Using financial calculator

Input:

PV= 5000

N= 10*2 = 20

I/Y = 10%/2 = 5%

Find PMT as 401.21

Dollar amount of each payment = $401.21

2: Composition of the two payments will be as under (Check workings)                        

Period

StartingBalance

Interest

Principal

EndingBalance

TotalInterest

1

$5,000.00

$250.00

$151.21

$4,848.79

$250.00

2

$4,848.79

$242.44

$158.77

$4,690.01

$492.44

Interest in first payment = $250

3: Principal = $151.21

4: Option 1 is true

(We see that the interest has reduced from 250 to 242.44 while principal repayment increased from 151.21 to 158.77)

WORKINGS

AutoSave Loan Amortization Excel (Product Activation Failed) Sign in File Insert Page Layout Formulas Data Review View Help Tell me what you want to do Home Share Calibri Wrap Text General Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- .. Sort & Find & Paste B , u . Ld 2.Д. 용 트經垣 Merge & Center· s-% , :08 g Filter Select Clipboard Alignment Number Styles Cells Editing АЗ 4 5 Loan Amount Interest Rate Term in Years Monthly Payment 6 5000 PMT(B6/2,(C6*2),-A6) Period StartingBalance Interest Principal B9*(BS6/2) B10 (BS6/2) DS6-C10 -B11 (BS6/2) DS6-C11 -812(BS6/2) =DS6-C12 B13 (BS6/2) DS6-C13 -814(BS6/2) =DS6-C14 B15 (BS6/2) DS6-C15 -B16 (BS6/2) DS6-C16 EndingBalance -B9-D9 -B10-D10 =B11-D11 B12-D12 B13-D13 -B14-D14 -B15-D15 -B16-D16 Totallnterest -C9 -F9+C10 -F10+C11 -F11+C12 DS6-C9 10 2 11 3 13 5 14 6 15 7 16 8 =E10 =E11 -E12 -E13 -E14 =E15 F12+C13 F13+C14 F14+C15 -F15+C16 Quarterly Semiannual MonthlyAnnual Ready + 100 C ENG 2:14 PM